HUBBLE SPACE TELESCOPE DAILY REPORT #986 PERIOD COVERED: 0800 EST Daily Status Report as of 278/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Uplinked Updated Ephemeris tables, Per MOSES 2871 1.2 Completed HSP/UV2 5005 (Gravitational Lenes - Part II _ Cycle 2 High) The High Speed Photometer will be used to determine a characteristic signature of variability in one component of the gravitational lens system KP1635+267. The time delay between the occurrences of the signature can be used to weigh the lens. The counts were very low for the onboard target acquisition, but the co-investigator said the data would be useful anyway. No other problems were reported. 1.3 Completed HSP/UV1 1091 (Ultraviolet Pulsations From X-Ray Pulsars) The High Speed Photometer was used to make single color photometric observations of the pulsating X-Ray (T-Tauri Class) binary star known as 1E1145.1-6. This is to find out if this object gives off ultraviolet emissions, as well as X-Ray pulsations. The observations were executed as scheduled, and no problems were reported. 1.4 Completed FGS 5054 (High Speed Astrometry - A Search for Planetary Companions to Low mass Stars) The Fine Guidance Sensors were used to make observations of the position of the star Proxima Centauri to determine the presence of a Jupiter sized planet in orbit about the target star. The OSS monitored the observations and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 1 At 278/0650Z the acquisition failed to acquire Fine Lock and went to Course Track. The FHST map at 0738 indicated errors of 16 arcseconds. Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 2 Successful : 0 FHST Failed to Update at 278/0911Z and 0914Z. Error box problems were indicated. 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete HSP/POL 4887 (X-Ray Binaries - Augmentation - Cycle 3 High) 3.2 Complete Two Sets of WFC 5023 (Calibration/ WFPC Cycle 3 Polarization Calibration) 3.3 Complete WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) 3.4 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.5 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.6 Complete FOC/96 3728 (Imaging Hot Stellar Content of Early Type Galaxies) 3.7 Complete WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.8 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 3.9 Complete FOC/96 5042 (FOC f/96 Aperture Location Monitoring) 3.10 Complete HSP/UV2 4943 (A Search for UV Variability in Nova Cyg 1992 - Cycle 2 High) 3.11 Complete HSP/POL 4942 (Active Galactic Nuclei - Cycle 2 High) 3.12 Complete GHRS/2 4067 (GHRS 1st Order Grating Spectral Lamp Test) 3.13 Complete PC 4764 (Investigation of the Dynamics of Binary Asteroids) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: JIS #4 (EVA #3) Simulation is scheduled for this coming Friday (Oct. 8, 1993). It will start at 1200Z and last for 12 hours.